Find the center of mass of a triangular lamina with vertices (0,0),(1,0) and (0,2)

if the density function is ρ(x,y)=1+3x+3y



1
Expert's answer
2022-03-01T04:12:27-0500

Let "A=(0,2), B=(1, 0), O=(0,0)."

Line OA: "x=0, 0\leq y\leq2"

LineAB: "y=-2x+2, 0\leq x\leq1"

Line OB: "y=0, 0\leq x\leq 1"


Given "\u03c1(x,y)=1+3x+3y"

Find the mass of the lamina


"m=\iint_D\rho(x,y)dA=\displaystyle\int_{0}^1\displaystyle\int_{0}^{2-2x}(1+3x+3y)dydx"




"=\displaystyle\int_{0}^1\big[y+3xy+{3y^2\over 2}\big]\begin{matrix}
 2-2x \\
 0 
\end{matrix}dx"




"=\displaystyle\int_{0}^1\big(2-2x+6x-6x^2+6-12x+6x^2\big)dx"


"=\big[-4x^2+8x\big]\begin{matrix}
 1 \\
 0
\end{matrix}=4(units\ of\ mass)"

Mass of the lamina is "4" units of mas.


Find the coordinates of the center of mass


"\bar{x}={1\over m}\iint_Dx\delta(x,y)dA"




"={1\over 4}\displaystyle\int_{0}^1\displaystyle\int_{0}^{1-x}x(1+3x+3y)dydx"

"={1\over 4}\displaystyle\int_{0}^1x\big[y+3xy+{3y^2\over 2}\big]\begin{matrix}
 2-2x \\
 0 
\end{matrix}dx"




"={1\over 4}\displaystyle\int_{0}^1\big(2x-2x^2+6x^2-6x^3\big)dx"

"+{1\over 4}\displaystyle\int_{0}^1\big(6x-12x^2+6x^3\big)dx"


"={1\over 4}\big[-{8\over 3}x^3+4x^2\big]\begin{matrix}
 1 \\
 0
\end{matrix}={1\over 3}"

"\bar{y}={1\over m}\iint_Dy\rho(x,y)dA"




"={1\over 4}\displaystyle\int_{0}^1\displaystyle\int_{0}^{1-x}y(1+3x+3y)dydx"




"={1\over 4}\displaystyle\int_{0}^1\big[{y^2\over 2}+{3xy^2\over 2}+y^3\big]\begin{matrix}
 2-2x \\
 0 
\end{matrix}dx"




"={1\over 4}\displaystyle\int_{0}^1\big(2-4x+2x^2+6x-12x^2+6x^3\big)dx"

"+{1\over 4}\displaystyle\int_{0}^1\big(8-24x+24x^2-8x^3\big)dx"

"={1\over 4}\big[-{1\over 2}x^4+{14\over 3}x^3-11x^2+10x\big]\begin{matrix}
 1 \\
 0
\end{matrix}={19\over 24}"

Center of mass "\big(\dfrac{1}{3},\dfrac{19}{24}\big)"